OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sin^2x + sin x = 0

OpenStudy (jtvatsim):

what have you tried so far? Usually these types of questions take a little algebra to set up.

OpenStudy (anonymous):

i havent been able to do much but sin x = 1/cscx

OpenStudy (jtvatsim):

ok, how about if you factor a sin(x) out of the left side? like this: \[\sin(x)\cdot(\sin(x)+1) =0\]

OpenStudy (anonymous):

ok what about the sin

OpenStudy (jtvatsim):

From algebra class we now know that: \[\sin(x) = 0 \enspace \text{or} \enspace \sin(x) + 1 = 0\]

OpenStudy (jtvatsim):

sorry, did you have a question about the sin? Which part?

OpenStudy (anonymous):

no i figured it out

OpenStudy (jtvatsim):

ok, do you think you can figure the rest out? :)

OpenStudy (anonymous):

graph sinx

OpenStudy (jtvatsim):

sure you can do that or if you know the unit circle you can use that too, just find the points where sin is 0 and where sin is -1 and you should be set.

OpenStudy (anonymous):

so pi/3 and 5pi/3 right

OpenStudy (jtvatsim):

not, quite, your graph of sin should be like this...

OpenStudy (jtvatsim):



OpenStudy (jtvatsim):

so we are supposed to look in the interval [0, 2pi) which means we should start at x = 0 and go up to 2pi, but don't use 2pi.

OpenStudy (jtvatsim):

so sin(0) = 0 -> x = 0 and sin(pi) = 0 -> x = pi for the -1 solutions sin(3pi/2) = -1 -> x = 3pi/2

OpenStudy (jtvatsim):

even though sin(2pi) does equal 0 , we ignore it since it is not in our interval

OpenStudy (anonymous):

oh okay thank you

OpenStudy (anonymous):

i actually have one more question

OpenStudy (anonymous):

Write the expression as the sine, cosine, or tangent of an angle. cos 96° cos 15° + sin 96° sin 15

OpenStudy (anonymous):

it would be cos81